OpenStudy (anonymous):

A and S are on circle D. 3434 Divided by 4

OpenStudy (anonymous):

\[area = \frac{ \theta r^{2} }{ 2 }\] make sure theta is in radians though another way is just to take a ratio like 118 degrees / 360 degrees = area of sector / area of circle so \[\frac{ 118 }{ 360 } = \frac{ sector }{ \pi r^{2} }\]

OpenStudy (anonymous):

so I divide that?

OpenStudy (anonymous):

yeah, take 118/360, then multiply that by pi r^2

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

i get d

OpenStudy (anonymous):

\[\frac{ 118 }{ 360 } = \frac{ x }{ \pi * 13^{2} }\]

OpenStudy (anonymous):

Ohhh ok I see now